# Perfect Wave Transfer in Continuous Quantum Systems

### v1

@@ 2. Example 2:
By comparing the terms appearing in Eqs. (42) and (40), it follows that $|F(t)|$ for certain $t$ equals $|\langle\Phi_{1}|\Phi_{1}\rangle|$ for all $\xi^{\pm}_{1}(x)$ if and only if $\underline{v}(x)$ is even, meaning $\underline{f}(x)$ must be odd. This in turn implies $\underline{f}(L)=L$ and $\underline{v}_{0}=v_{0}$ given by Eq. (7), which means that equality holds for all $t=(2m+1)L/v_{0}$ due the $2L$ periodicity of $G^{\pm}_{\Phi\Phi}(x)$. [Moreover, for $\xi^{\pm}_{2}(x)=\xi^{\mp}_{1}(-x)$, one has $|\langle\Phi_{1}|\Phi_{1}\rangle|=|\langle\Phi_{2}|\Phi_{2}\rangle|$ when $v(x)$ is even.] Since $\underline{v}(x)=v(x)$ for all $x\in[-L/2,L/2]$, this proves the claim that one-particle PWT is exhibited for all even inhomogeneous CFTs.
As a last remark, one can unfold in a different way to take the spin/statistics of the primary fields into account, so that, e.g., one gets anti-periodic BCs for fermions on the system of size $2L$, and twisted BCs for general anyonic fields. This can be used to obtain simple explicit formulas for the two-point correlations of any primary fields. However, since we are interested in recombinations of waves traveling in different directions, it is necessary to identify so as to always have periodic BCs. In any case, for our purposes, all we needed was the spacetime dependence of the correlations, which are fixed on general grounds in CFT, circumventing any need for explicit formulas.
@@
In this part, we first show that reflection symmetry is a necessary condition for PWT for any theory; here under the assumption that the Hamiltonian has pure point spectrum. Second, we give the details for how to use SL theory to diagonalize the inhomogeneous TLL Hamiltonian (8) and to formulate the PWT requirements on the spectrum. Third, we review some details on Weyl’s law applied to SL operators and, fourth, extend the previous discussion to TLL Hamiltonians with a mass term. Fifth, we comment on the Liouville normal form used to state the results for the inverse SL problem in the main text.
@@
Before considering TLL theory, we note that all models studied in the main text have pure point spectrum [as long as the integral defining $1/v_{0}$ in Eq. (7) converges]. Thus, we first consider a general Hamiltonian $H$ with pure point spectrum and show that for such a system to exhibit PST/PWT, it is necessary that $H$ commutes with the parity operator $\mathcal{P}$, which maps $x\mapsto-x$. In other words, we will show that reflection symmetry is a necessary condition.
To prove the statement, let $E_{n}$ denote the eigenenergies of $H$ with the corresponding energy eigenstates $|n\rangle$. On general grounds, the latter form a complete orthonormal basis for the Hilbert space $\mathcal{F}$ of the theory. Let $|\Phi_{1}\rangle$ be an arbitrary state in $\mathcal{F}$ and let $|\Phi_{2}\rangle\equiv\mathcal{P}|\Phi_{1}\rangle$. (We recall that $\mathcal{P}$ is a unitary operator satisfying $\mathcal{P}^{2}=1$, meaning that $\mathcal{P}^{-1}=\mathcal{P}^{\dagger}=\mathcal{P}$.) As usual, define $F(t)\equiv\langle\Phi_{2}|\mathrm{e}^{-\mathrm{i}Ht}|\Phi_{1}\rangle=\langle\Phi_{1}|\mathcal{P}\mathrm{e}^{-\mathrm{i}Ht}|\Phi_{1}\rangle$. For general PST/PWT, we require that there exists $T>0$ such that $|F(T)|=|\langle\Phi_{1}|\Phi_{1}\rangle|$ for any state $|\Phi_{1}\rangle$. In particular, this must hold for any energy eigenstate $|n\rangle$, implying $|\langle n|\mathcal{P}|n\rangle|=1$. If $|n\rangle|$ is not an eigenstate of $\mathcal{P}$, it is straightforward to show that this equality cannot hold. Consequently, all energy eigenstates must also be eigenstates of $\mathcal{P}$, which implies that $[\mathcal{P},H]=0$.
@@
We now turn to inhomogeneous TLLs studied in the main text. Inserting the expansions (14) into Eq. (10) yields
$H_{\textrm{iTLL}}=\frac{1}{2\pi}[\pi\Pi_{0}]^{2}+\frac{1}{2\pi}\sum_{n=1}^{\infty}\left.:\!\left([\pi\Pi_{n}]^{2}+\lambda_{n}\varphi_{n}^{2}\right)^{2}\!:\right.,$ (43)
where $\varphi_{0}$ is absent (when $E_{0}=0$), no different from the standard case with constant $K(x)$ [and $v(x)$]. For $n\geq 1$, let
@@
in terms of bosonic creation and annihilation operators $a^{\dagger}_{n}$ and $a^{\vphantom{\dagger}}_{n}$, satisfying $\bigl{[}a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr{]}=\delta_{n,n^{\prime}}$, as in the main text. Inserting Eq. (44) into Eq. (43) yields Eqs. (15)–(16).
@@
where we omitted the zero modes (as they have to be treated separately when $E_{0}=0$) and used $\mathrm{e}^{\mathrm{i}H_{\textrm{iTLL}}t}a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}H_{\textrm{iTLL}}t}=a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}$ and $\mathrm{e}^{\mathrm{i}H_{\textrm{iTLL}}t}a^{\dagger}_{n}\mathrm{e}^{-\mathrm{i}H_{\textrm{iTLL}}t}=a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}$. Generally, any observable $\mathcal{O}(x,t)$ in this algebra will contain contributions of the form $a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}u_{n}(x)$ and $a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}u_{n}(x)$. Now, in order for PWT to be exhibited, Eq. (1) must hold for any state, which means that it holds true without expectation values, i.e., $\mathcal{O}(x,0)$ must equal $\mathcal{O}(-x,T)$ at some later time $T>0$ as an operator identity. Taking $\mathcal{O}=\varphi$ and using the linear independence of the generators $a^{\vphantom{\dagger}}_{n}$ and $a^{\dagger}_{n}$, Eq. (45) yields
@@
for a common $T>0$, where we included the trivial case $n=0$ [recall that $u_{0}(x)=\textrm{const}$ and $E_{0}=0$ when $q(x)=0$ in Eq. (11)]. Since $\mathcal{P}u_{n}(x)=u_{n}(-x)$, Eq. (46) implies that each $u_{n}(x)$ must be an eigenfunction of the parity operator $\mathcal{P}$, whose eigenvalues are $+1$ or $-1$ (since $\mathcal{P}^{2}=1$). It follows that all eigenfunctions must have definite parity, and the only possibility for PWT to hold is that the corresponding eigenenergies satisfy $E_{n}T=\pi m_{n}$ for integers $m_{n}$ for each $n$, with $m_{n}$ even (odd) if $u_{n}(x)$ has even (odd) parity.
Finally, it is straightforward to show the following statement: The eigenfunctions $u_{n}(x)$ of a regular SL operator $\mathcal{A}$ have definite parity if and only if $[\mathcal{A},\mathcal{P}]=0$. For our operator in Eqs. (11)–(12), this means that both $v(x)$ and $K(x)$ must be even functions (using that they must be positive to exclude the possibility that both are instead odd).
Given the above, together with Properties (i)(iv) for regular SL operators stated in the main text, we have derived our main result on the spectral requirements for an inhomogeneous TLL to exhibit PWT.
Moreover, as stated in the main text, reversing the logic, namely, given a spectrum that satisfy the spectral properties for PWT to be exhibited, our formulation of PWT for inhomogeneous TLLs as an inverse SL problem follows.
@@
Weyl’s law applied to SL operators, see, e.g., [78, 79], states that the number $N(\lambda)$ of eigenvalues below a given $\lambda>0$ asymptotically behaves as
@@
for large $\lambda$. This holds under certain technical conditions on the coefficients $w(x)$, $p(x)$ in Eq. (11) [which are satisfied assuming our $v(x)$, $K(x)$ are positive and smooth functions]. Using Eqs. (12) and (16) as well as Property (i) in the main text, it follows that
@@
for large $n$, where in the last step we used Eq. (7). Thus, assuming that the integral defining $1/v_{0}$ converges, then this gives the stated asymptotic behavior, $E_{n}=\pi v_{0}n/L+o(n)$, in the main text. (Another possibility would be if the integral diverges, in which case $E_{n}\sim\textrm{const}\times n^{\gamma}$ for $\gamma<1$, but this is inconsistent with that the eigenenergies must be proportional to integers $m_{n}$ with a common proportionality constant in order for the theory to exhibit PWT.)
@@
We now extend to SL operators (11) with $q(x)\neq 0$, corresponding to an inhomogeneous (possibly imaginary) mass $M(x)=\sqrt{-q(x)/v_{0}^{2}v(x)}$ by adding $v_{0}^{2}M(x)^{2}\varphi(x)^{2}$ inside the big parentheses in Eq. (8):
$H_{\textrm{iTLL}}\equiv\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}v(x)\!\left.:\!\biggl{(}\frac{[\pi\Pi(x)]^{2}}{K(x)}+K(x)[\partial_{x}\varphi(x)]^{2}+v_{0}^{2}M(x)^{2}\varphi(x)^{2}\biggr{)}\!:\right.\!,$ (49)
Indeed, it is straightforward to extend the previous discussion to such models. For instance, the PWT requirements on the spectrum are the same, except the spectrum does not start at zero: For $q(x)\neq 0$, a constant $u_{0}(x)$ no longer solves Eq. (13), thus $E_{0}>0$. Instead, we have a “massive” inhomogeneous TLL, for which one can show:
A “massive” inhomogeneous TLL exhibits PWT if and only if $v(x)$, $K(x)$, $q(x)$ are even
and all eigenvalues are of the form $E_{n}=\pi\tilde{m}_{n}/T$ for certain integers $\tilde{m}_{n}$.
Here, these integers $\tilde{m}_{n}=m_{n}+c$ in terms of the integers $m_{n}$ for the massless case in the main text, where we allow for an overall shift by a finite even integer $c$. [The latter must be even so that $\tilde{m}_{n}$ is even (odd) if $n$ is even (odd), as required due to Property (ii) in the main text.] Moreover, given such a spectrum, the results in [82, 83] still imply that $\hat{q}(y)$ is uniquely determined by the spectrum, but this can no longer uniquely determine $K(x)$ since the first term in Eq. (18) is now non-zero in general.
As an example, we revisit the case with $v(x)=v\sqrt{1-(2x/L)^{2}}$ and $K(x)=K\bigl{[}1-(2x/L)^{2}\bigr{]}^{\alpha}$ for $\alpha>-1/2$ and constants $v,K>0$ encountered in the main text, whose eigenfunctions were given by Gegenbauer polynomials $C^{(\alpha)}_{n}(2x/L)$, but now we include a fine-tuned potential $q(x)=-(2v\alpha/L)^{2}w(x)$, where we recall that $w(x)=K(x)/v(x)$. This fine-tuned potential allows us to complete the square so that the eigenvalues in Eq. (13) are $\lambda_{n}=(\pi v_{0}/L)(n+\alpha)^{2}$ with $v_{0}=2v/\pi$, meaning that $E_{n}=\pi(n+\alpha)v_{0}/L$. Consequently, any $c=\alpha=0,2,4,\ldots$ works, yielding a discrete family of models with fine-tuned inhomogeneous mass that exhibit PWT. We stress that the mass in this example satisfies $M(x)\propto 1/L$ and therefore vanishes in the thermodynamic limit. Thus, one should not expect hallmarks of massive theories, such as exponential decay of two-point correlations with respect to the spatial separation.

### v3

@@ 2. Example 2:
By comparing the terms appearing in Eqs. (42) and (40), it follows that $|F(t)|$ for certain $t$ equals $|\langle\Phi_{1}|\Phi_{1}\rangle|$ for all $\xi^{\pm}_{1}(x)$ if and only if $\underline{v}(x)$ is even, meaning $\underline{f}(x)$ must be odd. This in turn implies that $\underline{f}(L)=L$ and $\underline{v}_{0}=v_{0}$ given by Eq. (7), which means that equality holds for all $t=(2m+1)L/v_{0}$ due the $2L$ periodicity of $G^{\pm}_{\Phi\Phi}(x)$. [Moreover, for $\xi^{\pm}_{2}(x)=\xi^{\mp}_{1}(-x)$, one has $|\langle\Phi_{1}|\Phi_{1}\rangle|=|\langle\Phi_{2}|\Phi_{2}\rangle|$ when $v(x)$ is even.] Since $\underline{v}(x)=v(x)$ for all $x\in[-L/2,L/2]$, this proves the claim that one-particle PWT is exhibited for all even inhomogeneous CFTs.
As a last remark, it is possible to unfold in different ways to incorporate the statistics of the primary fields. E.g., so that, for the system of length $2L$, one gets antiperiodic BCs for fermions and twisted BCs for general anyonic fields. However, since we are interested in recombinations of waves traveling in different directions, we found it necessary to identify right and left movers so as to always have periodic BCs.
@@
In this part, we first show that reflection symmetry is a necessary condition for PWT in any system; here, under the assumption that the Hamiltonian has pure point spectrum. Second, we give the details for how to use SL theory to diagonalize the inhomogeneous TLL Hamiltonian (8) and how to formulate the PWT requirements on the spectrum. Third and fourth, respectively, we review certain details for Weyl’s law and the Bohr-Sommerfeld/WKB quantization condition applied to SL operators, and, fifth, extend the previous discussion to TLL Hamiltonians with a mass term. Sixth, we comment on the Liouville normal form used to state the results for the inverse SL problem in the main text.
@@
Before considering TLL theory, we note that all models studied in the main text have pure point spectrum [as long as the integral defining $1/v_{0}$ in Eq. (7) converges]. Thus, we first consider a general Hamiltonian $H$ with pure point spectrum and show that for such a system to exhibit PWT or PST, it is necessary that $H$ commutes with the parity operator $\mathcal{P}$, which maps $x\mapsto-x$. In other words, we will show that reflection symmetry is a necessary condition.
To prove the statement, let $E_{n}$ denote the eigenenergies of $H$ with the corresponding energy eigenstates $|n\rangle$. On general grounds, the latter form a complete orthonormal basis for the Hilbert space $\mathcal{F}$ of the theory. Let $|\Phi_{1}\rangle$ be an arbitrary state in $\mathcal{F}$ and let $|\Phi_{2}\rangle\equiv\mathcal{P}|\Phi_{1}\rangle$. (We recall that $\mathcal{P}$ is a unitary operator satisfying $\mathcal{P}^{2}=1$, meaning that $\mathcal{P}^{-1}=\mathcal{P}^{\dagger}=\mathcal{P}$.) As usual, define $F(t)\equiv\langle\Phi_{2}|\mathrm{e}^{-\mathrm{i}Ht}|\Phi_{1}\rangle=\langle\Phi_{1}|\mathcal{P}\mathrm{e}^{-\mathrm{i}Ht}|\Phi_{1}\rangle$. Full PWT, as defined in Eq. (1), requires that there exists a common shortest possible time $T>0$ such that $|F(T)|=|\langle\Phi_{1}|\Phi_{1}\rangle|$ for all states $|\Phi_{1}\rangle$. In particular, this must hold for any energy eigenstate $|n\rangle$, implying that $|\langle n|\mathcal{P}|n\rangle|=1$. If $|n\rangle$ is not an eigenstate of $\mathcal{P}$, then it is straightforward to show that this equality cannot hold. Consequently, all energy eigenstates must also be eigenstates of $\mathcal{P}$, which implies that $\mathcal{P},H]=0$.
@@
We now turn to inhomogeneous TLL theory studied in the main text. Inserting the expansions in Eq. (14) into the Hamiltonian (10) and using Eq. (13) yield
$H_{\textrm{iTLL}}=\frac{1}{2\pi}[\pi\Pi_{0}]^{2}+\frac{1}{2\pi}\sum_{n=1}^{\infty}\left.:\!\hskip-0.5pt\Bigl([\pi\Pi_{n}]^{2}+\lambda_{n}\varphi_{n}^{2}\Bigr)\hskip-0.5pt\!:\right.$. (43)
Here, $\varphi_{0}$ is absent (when $E_{0}=0$), no different from the standard case with constant $K(x)$ [and $v(x)$]. For $n\geq 1$, let
@@
in terms of bosonic creation and annihilation operators $a^{\dagger}_{n}$ and $a^{\vphantom{\dagger}}_{n}$ satisfying $\bigl[a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr]=\delta_{n,n^{\prime}}$ and $a_{n}|\Omega\rangle=0$, where $|\Omega\rangle$ denotes the ground state of $H_{\textrm{iTLL}}$; cf. Sec. 3.2 in [37]. Inserting Eq. (44) into Eq. (43) yields Eqs. (15)–(16).
@@
where we omitted the zero modes (they must be treated separately when $E_{0}=0$, but in that case they will have no consequence for our discussion of PWT) and used $\mathrm{e}^{\mathrm{i}H_{\textrm{iTLL}}t}a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}H_{\textrm{iTLL}}t}=a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}$ and $\mathrm{e}^{\mathrm{i}H_{\textrm{iTLL}}t}a^{\dagger}_{n}\mathrm{e}^{-\mathrm{i}H_{\textrm{iTLL}}t}=a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}$. Generally, any observable $\mathcal{O}(x,t)$ in this algebra will contain contributions of the form $a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}u_{n}(x)$ and $a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}u_{n}(x)$. Now, in order for PWT to be exhibited, Eq. (1) must hold for any state in the Hilbert space, which means that it must hold without taking expectation values. I.e., $\mathcal{O}(x,0)$ must equal $\mathcal{O}(-x,T)$ at some later time $T>0$ as an operator identity. Taking $\mathcal{O}=\varphi$ and using the linear independence of the generators $a^{\dagger}_{n}$ and $a^{\vphantom{\dagger}}_{n}$, Eq. (45) implies that
@@
for a common shortest possible $T>0$, where we included the trivial case $n=0$ [recall that $u_{0}(x)=\textrm{const}$ and $E_{0}=0$ when $q(x)=0$ in Eq. (11)]. Since $\mathcal{P}u_{n}(x)=u_{n}(-x)$, Eq. (46) implies that each $u_{n}(x)$ must be an eigenfunction of the parity operator $\mathcal{P}$, whose eigenvalues are +1+1 or 1-1 (since $\mathcal{P}^{2}=1$). It follows that all eigenfunctions must have definite parity, and the only possibility for PWT to hold is that the corresponding eigenenergies satisfy $E_{n}T=\pi m_{n}$ for integers $m_{n}$, with $m_{n}$ even (odd) if $u_{n}(x)$ has even (odd) parity.
Finally, it is straightforward to show the following statement: The eigenfunctions $u_{n}(x)$ of a regular SL operator $\mathcal{A}$ have definite parity if and only if $\mathcal{A},\mathcal{P}]=0$. For our operator in Eqs. (11)–(12), this means that both $v(x)$ and $K(x)$ must be even functions (using that they must be positive to exclude the possibility that both are instead odd).
Given the above, together with Properties (i)(iv) for regular SL operators stated in the main text, we have derived our statement in terms of spectral requirements for an inhomogeneous TLL to exhibit PWT.
Moreover, as stated in the main text, reversing the logic, i.e., assuming we are given a spectrum that satisfies the spectral requirements for PWT, we obtain our formulation of PWT for inhomogeneous TLLs as an inverse problem.
@@
Weyl’s law applied to SL operators of the general form (11) states that the number $N(\lambda)$ of eigenvalues below a given $\lambda>0$ asymptotically behaves as
@@
for large $\lambda$; see, e.g., [97, 98]. This holds under certain technical conditions on the functions $w(x)$, $p(x)$, and $q(x)$ in Eq. (11), which are satisfied assuming our $v(x)$ and $K(x)$ are positive and smooth functions [if $q(x)$ is nonzero, then it is also assumed to be smooth]. Using Eqs. (12) and (16) as well as Property (i) in the main text, it follows that
@@
for large $n$, where in the last step we used Eq. (7). Thus, assuming that the integral defining $1/v_{0}$ converges, this gives the stated asymptotic behavior, $E_{n}=\pi v_{0}n/L+o(n)$, in the main text. (Another possibility would be if the integral diverges, in which case $E_{n}\sim\textrm{const}\times n^{\gamma}$ for $\gamma<1$, but this is inconsistent with that the eigenenergies must be proportional to integers $m_{n}$ with a common proportionality constant, in order for the theory to exhibit PWT.)
Exact Bohr-Sommerfeld/WKB quantization condition
A stronger version of Weyl’s law for our class of operators can be deduced from an exact generalization of the usual formulation of the Bohr-Sommerfeld/WKB quantization condition. To state this condition, we consider our Sturm-Liouville problem written on Liouville normal form (19):
$\bigl[-\partial_{y}^{2}+V(y)\bigr]\mathfrak{u}(y)=\Lambda\mathfrak{u}(y)$ (49)
on the interval $I\equiv[-L/2,L/2]\ni y$, where we introduced $V(y)\equiv\hat{q}(y)/v_{0}^{2}$ and $\Lambda\equiv\lambda/v_{0}^{2}$ and recall that we impose Neumann BCs. The first terms of the exact quantization condition for the $n$th eigenvalue $\Lambda_{n}$ can then be stated as
$\phi(\Lambda_{n})=\pi n$ (50)
with
$\phi(\Lambda)=\phi_{0}(\Lambda)+\phi_{1}(\Lambda)+O(\Lambda^{-5/2}),\vskip-7.11317pt$ (51)
$\phi_{0}(\Lambda)\equiv\int_{I}\mathrm{d}y\,\sqrt{\Lambda-V(y)},\qquad\phi_{1}(\Lambda)\equiv-\frac{1}{4}\frac{\mathrm{d}}{\mathrm{d}\Lambda}\int_{I}\mathrm{d}y\,\frac{V^{\prime\prime}(y)}{\sqrt{\Lambda-V(y)}}.$ (52)
This follows from [110, 111, 112] with the difference that we have no soft turning points but instead a finite interval with Neumann BCs: There is thus no Maslov index, and so the right-hand side of Eq. (50) is $\pi n$ instead of $\pi(n+1/2)$; cf. Sec. 2.4 in [113].
First, we consider $\phi_{1}(\Lambda)$. Using the expansion $1/\sqrt{\Lambda-V(y)}=\Lambda^{-1/2}\bigl[1+V(y)/2\Lambda+O(\Lambda^{-2})\bigr]$ for large $\Lambda$, we obtain
$\int_{I}\mathrm{d}y\,\frac{V^{\prime\prime}(y)}{\sqrt{\Lambda-V(y)}}=\frac{1}{\sqrt{\Lambda}}\int_{I}\mathrm{d}y\,V^{\prime\prime}(y)+O(\Lambda^{-3/2}),$ (53)
which implies that
$\phi_{1}(\Lambda)=\frac{L}{8\Lambda^{3/2}}\Delta+O(\Lambda^{-5/2})$ (54)
with
$\Delta\equiv\frac{V^{\prime}(L/2)-V^{\prime}(-L/2)}{L}.$ (55)
Recall that $V(y)$ must be an even function on $I$ (for the system to exhibit PWT) and thus $V^{\prime}(y)$ will be odd. Second, we consider $\phi_{0}(\Lambda)$. Using the expansion $\sqrt{\Lambda-V(y)}=\Lambda^{1/2}\bigl[1-V(y)/2\Lambda-V(y)^{2}/8\Lambda^{2}+O(\Lambda^{-3})\bigr]$ for large $\Lambda$, we find
$\phi_{0}(\Lambda)=\sqrt{\Lambda}L\left(1-\frac{1}{2\Lambda}a_{1}-\frac{1}{8\Lambda^{2}}a_{2}\right)+O(\Lambda^{-5/2})$ (56)
with
$a_{1}\equiv\frac{1}{L}\int_{I}\mathrm{d}y\,V(y),\qquad a_{2}\equiv\frac{1}{L}\int_{I}\mathrm{d}y\,V(y)^{2},$ (57)
which can be interpreted as the first two moments of the potential $V(y)$.
Inserting the above into Eqs. (50)–(52) yields
$\sqrt{\Lambda_{n}}\left(1-\frac{1}{2\Lambda_{n}}a_{1}-\frac{1}{8\Lambda_{n}^{2}}[a_{2}-\Delta]+O(\Lambda_{n}^{-3})\right)=\frac{\pi n}{L}.$ (58)
Since Eq. (16) implies $\Lambda_{n}=\lambda_{n}/v_{0}^{2}=(E_{n}/v_{0})^{2}$ and since $E_{n}=\pi m_{n}/T$ for integers $m_{n}$ satisfying the conditions (a)(b) on page (a) of the main text, the quantization condition becomes
$m_{n}\left(1-\frac{L^{2}}{2\pi^{2}m_{n}^{2}}a_{1}-\frac{L^{4}}{8\pi^{4}m_{n}^{4}}[a_{2}-\Delta]+O(m_{n}^{-6})\right)=n,$ (59)
where we used that $v_{0}=L/T$ by definition of $T$ as the common shortest possible time (for all modes) for the system to exhibit PWT. Considering the difference $m_{n+1}-m_{n}$, it follows from Eq. (59) that $m_{n+1}-m_{n}=1$ for all $n>n_{0}$ for some sufficiently large positive integer $n_{0}$. The solution of this recursion formula is $m_{n}=n+a$ for some $a\in\mathbb{R}$ for $n>n_{0}$. In other words, $m_{n}=n\bigl[1+O(n^{-1})\bigr]$ for large $n$, which compared with Eq. (59) implies
$a_{1}=0,\qquad a_{2}-\Delta=0.$ (60)
I.e., the potential $V(y)$ must have a vanishing mean, which implies that $a_{2}$ is the variance, which must equal $\Delta$. [Note that Eqs. (55) and (57) then imply that $V^{\prime}(L/2)\geq 0$, since $\Delta=a_{2}\geq 0$ could otherwise not hold.]
The interval $I$ with Neumann BCs can be unfolded to [L/2,3L/2][-L/2,3L/2] of length $2L$ with periodic BCs (similar to Part A), which in turn can be viewed as the circle $S^{1}_{2L}$. If we impose that $V(y)$ must be regular not only before but also after this unfolding procedure, then the only possibility is $V^{\prime}(L/2)=0$, since $V(y)$ must be even, from which it follows that $\Delta=0$ in Eq. (55). Consequently, both the mean $a_{1}$ and the variance $a_{2}$ of the potential $V(y)$ vanish in that case, which then implies that $\hat{q}(y)=v_{0}^{2}V(y)$ must be zero (almost everywhere). From Eq. (18) with $q(x)=0$ and reflection symmetry, it follows that $K(x)$ must be constant for PWT to be exhibited in an inhomogeneous TLL with the assumption of regularity after unfolding to $S^{1}_{2L}$. However, if such regularity is not assumed, then the quantization condition (59) does not prevent $\hat{q}(y)=v_{0}^{2}V(y)$ from being nonzero, which in turn corresponds to a nonconstant $K(x)$.
@@
We now extend to SL operators (11) with a smooth $q(x)\neq 0$, corresponding to an inhomogeneous (possibly imaginary) mass $M(x)=\sqrt{-q(x)/v_{0}^{2}v(x)}$ by adding $v_{0}^{2}M(x)^{2}\varphi(x)^{2}$ inside the big parentheses in Eq. (8):
$H_{\textrm{miTLL}}\equiv\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}v(x)\!\left.:\!\hskip-0.5pt\biggl(\frac{[\pi\Pi(x)]^{2}}{K(x)}+K(x)[\partial_{x}\varphi(x)]^{2}+v_{0}^{2}M(x)^{2}\varphi(x)^{2}\biggr)\hskip-0.5pt\!:\right.\!$. (61)
Indeed, it is straightforward to extend the previous discussion to such models. For instance, the PWT requirements on the spectrum are the same, except the spectrum does not start at zero: For $q(x)\neq 0$, a constant $u_{0}(x)$ no longer solves Eq. (13), thus $E_{0}>0$. Instead, we have an inhomogeneous TLL with a mass term, for which one can show the following statement: “Massive” inhomogeneous TLLs exhibit PWT as defined in Eq. (1) if and only if $v(x)$, $K(x)$, and $q(x)$ are even and the eigenenergies are $E_{n}=\pi\tilde{m}_{n}/T$ for integers $\tilde{m}_{n}$ so that
  1. (a)
    $0\leq\tilde{m}_{0}<\tilde{m}_{1}<\ldots$ with each $\tilde{m}_{n}$ even or odd if $n$ is even or odd,
  2. (b)
    $\tilde{m}_{n}\sim n$ for large $n$, i.e., $\tilde{m}_{n}=n+o(n)$.
The integers in the result above can be written as $\tilde{m}_{n}=m_{n}+k$ in terms of the integers $m_{n}$ for the massless case in the main text, allowing for an overall shift by a finite even integer $k$. [The latter must be even so that (a) above is satisfied, as required due to Property (ii) in the main text.] Now, given such a spectrum, the results in [105, 106] still imply that $\hat{q}(y)$ is uniquely determined by it, but this can no longer uniquely determine $K(x)$ since the first term on the right-hand side of Eq. (18) is nonzero in general.
As an example, we revisit the case with $v(x)=v\sqrt{1-(2x/L)^{2}}$ and $K(x)=K\bigl[1-(2x/L)^{2}\bigr]^{\alpha}$ for $\alpha>-1/2$ and constants $v,K>0$ encountered in the main text, whose eigenfunctions were given by Gegenbauer polynomials $C^{(\alpha)}_{n}(2x/L)$, but now we include a fine-tuned potential $q(x)=-(2v\alpha/L)^{2}w(x)$, where we recall that $w(x)=K(x)/v(x)$. This fine-tuned potential allows us to complete the square so that the eigenvalues in Eq. (13) are $\lambda_{n}=(\pi v_{0}/L)(n+\alpha)^{2}$ with $v_{0}=2v/\pi$, meaning that $E_{n}=\pi(n+\alpha)v_{0}/L$. Consequently, any $\alpha=k=0,2,4,\ldots$ works, yielding a discrete family of models with fine-tuned inhomogeneous mass that exhibit PWT. We stress that the mass in this example satisfies $M(x)\propto 1/L$ and therefore vanishes in the thermodynamic limit. Thus, one should not expect hallmarks of massive theories, such as exponential decay of two-point correlations with respect to the spatial separation.
From the theory of regular SL operators and the Bohr-Sommerfeld/WKB quantization condition in the previous section, if one assumes regularity also after unfolding to $S^{1}_{2L}$, then “massive” inhomogeneous TLLs must lie in the orbit of $\hat{q}(y)=0$ in order to exhibit PWT. I.e., in that case, while $K(x)$ does not have to be constant, the choice of $q(x)$ must be such that $\hat{q}(y)$ given by Eq. (18) is still zero, corresponding to an exactly linear spectrum as stated in the main text. However, the above Gegenbauer example is irregular, meaning that it does not have to lie in the orbit of $\hat{q}(y)=0$ (which it does not for $\alpha=2,4,\ldots$) while exhibiting PWT.